OpenStudy (sh3lsh):

Multivariable question! Find the equation of the plane tangent to the graph of the function z(x,y)=2x^2-7y^2-x-6y+3 at the point (x,y)=(5,-1). I say that plane z = 19x+8y-87, but it's supposedly incorrect, help?

OpenStudy (loser66):

Let see: \(f'_x= 4x-1\) hence at (5, -1) , \(f'_x= 19\) \(f'_y= -14y -6\), at (5,-1), \(f'_y= 8\) hence, the tangent line equation is \(Z= 19(x-5) +8(y+1) = 19x-95+8y+8=19x+8y-87\)

OpenStudy (loser66):

Hence, we are the same boat. hehehe

OpenStudy (sh3lsh):

Haha, oh well. Thanks for confirming it!